OpenStudy (anonymous):

is this correct? 1/x^2 dx

OpenStudy (anonymous):

I got x^-1 + C

OpenStudy (anonymous):

or is it -x^-1 + C

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ty :))

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

is it because i moved x^2 to the top?

OpenStudy (anonymous):

\[ \int {1\over x^2}dx=\int x^{-2}dx \]

OpenStudy (anonymous):

so the answer is x^-1 + C?

OpenStudy (anonymous):

\[={x^{-1}\over -1}+c\]

OpenStudy (anonymous):

ohh ok ty :)

OpenStudy (agent0smith):

Remember that after increasing the power by one (to -1), you need to divide by the new power (-1).

OpenStudy (agent0smith):

\[\Large \int\limits x^{-2}dx = - \frac{ 1 }{ x } +C\]

OpenStudy (anonymous):

yeah i forgot about that rule lol